# Lattice QCD Benchmark of Proton Helicity and Flavor-Dependent Unpolarized TMDPDFs at Physical Quark Masses

### v1

@@ IV Results
The bare matrix elements require renormalization. For CG quasi-TMD beam functions, this reduces to a multiplicative constant independent of the quark separation $\{b_{T},b_{z}\}$ and external hadron states Gao et al. (2024a). Following Refs. Bollweg et al. (2024); Avkhadiev et al. (2024), we apply the ratio
[…]
In Fig. 2, we present the real parts of renormalized matrix elements for isovector unpolarized ($\tilde{h}^{R,u_{v}-d_{v}}_{f_{1}}$), isoscalar unpolarized ($\tilde{h}^{R,u_{v}+d_{v}}_{f_{1}}$), and isovector helicity ($\tilde{h}^{R,\Delta u_{+}-\Delta d_{+}}_{g_{1L}}$) quasi-TMD beam functions at several $b_{T}$. This work focuses exclusively on the real part of the matrix elements. For unpolarized TMDPDFs, it corresponds to the valence-quark contribution $q_{v}$, while for helicity TMDPDFs, it includes both quark and antiquark contributions $\Delta q_{+}$. For the $x$ region considered in this work, the antiquark contributions are typically small and generally consistent with zero within statistical uncertainties Gao et al. (2024b, 2023).
Owing to the finite size of the hadron at realistic boosts, spatial correlators decay exponentially at large separations Gao et al. (2022). As shown in Fig. 2, $\tilde{h}^{R}$ becomes consistent with zero for $b_{z}\gtrsim 1$ fm within uncertainties. While the central values are already close to zero, the uncertainties in this region can introduce minor instabilities in the Fourier transform. To mitigate this and ensure numerical stability, we smoothly model the large-$b_{z}$ behavior by fitting the data in the range $b_{z}\in[b_{z}^{L},b_{z}^{L}+4a]$ using an exponential form $Ae^{-mb_{z}}$, constrained by a prior $m>0.1$GeV. As shown by the bands in Fig. 2, the extrapolated curves are consistent with the data and smoothly approach zero. It is important to note that the extrapolation has negligible impact on the final results. Even without it, a direct discretized Fourier transform (DFT) over the available data yields results that are nearly identical. The extrapolation primarily serves to suppress unphysical noise at large $b_{z}$, and its contribution is minimal. We further quantify the systematic uncertainty from the extrapolation by varying the lower bound $b_{z}^{L}$ between $10a$ and $16a$, and find the resulting variations in the quasi-TMD beam functions to be insignificant. To ensure a smooth and well-controlled reconstruction in $x$-space, we therefore perform a discretized Fourier transform over $b_{z}\in[0,b_{z}^{L}]$ and supplement it with an analytic continuation of the fitted exponential behavior over $b_{z}\in[b_{z}^{L},\infty]$.
@@
Figure 3: Lattice QCD results (filled symbols with error bars) for the ratio $R^{u-d}_{g_{1L}/f_{1}}(x,b_{T})$ (see Eq. (3)) of isovector helicity to isovector unpolarized TMDPDFs, shown as a function of $b_{T}$ at several values of $x$. Hatched bands indicate the corresponding phenomenological parameterization (MAP-Heli) from Ref. Bacchetta et al. (2024b).
@@
Figure 4: Lattice QCD results (filled symbols with error bars) for the ratio $R_{f_{1}}^{u/d}(x,b_{T})$ (see Eq. (4)) of up- to down-quark unpolarized TMDPDFs, shown as a function of $b_{T}$ at several values of $x$. Hatched bands denote corresponding phenomenological parameterizations from Ref. Bacchetta et al. (2024a) (MAP24) and Ref. Moos et al. (2025) (ART25).
After performing a Fourier transform over $b_{z}$, we obtain the helicity quasi-TMD beam function $\tilde{g}_{1L}^{\Delta u_{+}-\Delta d_{+}}(x,b_{T},P_{z})$ and the unpolarized quasi-TMD beam function $\tilde{f}_{1}^{u_{v}-d_{v}}(x,b_{T},P_{z})$ and $\tilde{f}_{1}^{u_{v}+d_{v}}(x,b_{T},P_{z})$. As discussed before, the ratios of these quasi-TMD beam functions are both factorization and renormalization scale independent, and correspond to the ratio of physical TMDPDFs up to power corrections.
Fig. 3 shows the RGI ratio $R^{u-d}_{g_{1L}/f_{1}}(x,b_{T})$, defined in Eq. (3), for several values of $x$. The $b_{T}$ dependence is weak, consistent with the flavor-independent global fit MAP-Heli Bacchetta et al. (2024b), particularly at $x=0.2$.
Fig. 4 shows the ratio $R_{f_{1}}^{u/d}(x,b_{T})$, defined in Eq. (4), for several $x$. For comparison, results from the flavor-dependent global fits MAP24 Bacchetta et al. (2024a) and ART25 Moos et al. (2025) are shown as bands. These exhibit differing $b_{T}$ trends.
Appendix C presents comparisons between CG and gauge-invariant formalisms. While CG results remain stable up to $b_{T}\approx 1$ fm, the signal in the gauge-invariant case degrades beyond $b_{T}\approx 0.4$ fm. Up to $b_{T}\approx 0.4$ fm, both methods yield consistent results, suggesting power corrections are subdominant at this level of precision.
These results provide valuable insight into the proton’s transverse structure. The flat $b_{T}$ dependence of $R^{u-d}_{g_{1L}/f_{1}}$ indicates that the transverse momentum distributions of helicity and unpolarized TMDPDFs are remarkably similar in the isovector channel. This suggests that helicity-dependent and helicity-independent transverse dynamics are not strongly differentiated at moderate $x$, and that longitudinal spin polarization has limited impact on the intrinsic transverse momentum profile. Such behavior offers a nonperturbative cross-check of phenomenological fits and is particularly important given the inconsistencies among current global extractions of the helicity TMDPDF. For the unpolarized case, the mild $b_{T}$ variation in $R_{f_{1}}^{u/d}$ points to a weak flavor dependence in the intrinsic transverse momentum distribution, implying that the primary differences between up and down quarks lie in their longitudinal structure. These findings are especially relevant in light of the tensions among existing global fits, which often rely on model-dependent assumptions in poorly constrained regions. Our lattice QCD results thus provide a robust first-principles benchmark for guiding future global analyses and reducing associated uncertainties.
V Conclusions
We presented the first lattice QCD results for the isovector helicity TMDPDF and the flavor-dependent unpolarized TMDPDFs for up and down quarks in the proton. Novel aspects include computations with physical up, down, and strange quark masses, discretizations preserving exact chiral symmetry at finite lattice spacing, and employing Coulomb-gauge-fixed quasi-TMD beam functions to reliably access nonperturbatively large transverse distances of about one fermi.
By constructing renormalization- and factorization-scale independent ratios, we achieve direct comparison with phenomenological extractions. This ratio eliminates uncertainties arising from scheme-dependent soft factors. Our results show a close similarity between the helicity and unpolarized isovector TMDPDFs at moderate $x$, and only mild flavor dependence in the unpolarized sector. These findings not only corroborate trends seen in recent global fits but also provide valuable nonperturbative input where phenomenology remains unconstrained or inconsistent.
While this study is based on a single lattice ensemble and does not yet include disconnected contributions, several factors support the robustness of our findings. The use of chiral fermions at physical quark masses, the consistency across multiple fit strategies, and the agreement between Coulomb-gauge-fixed and gauge-invariant results at small $b_{T}$ all indicate that other systematics might not be significant compared to the present statistical uncertainties. Moreover, previous lattice studies of PDFs and TMDs suggest that disconnected contributions and discretization effects are typically mild at current stage Alexandrou et al. (2021); Avkhadiev et al. (2024); Hua et al. (2022); Gao et al. (2022). Taken together, these considerations suggest that the key features observed in our results—such as the similarity between helicity and unpolarized TMDPDFs and the weak flavor dependence in transverse momentum genuinely reflects proton’s nonperturbative structures. This work establishes a strong foundation for future precision lattice investigations with improved control of systematics and provides timely input for global analyses.
Acknowledgements
@@
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704, Contract No. DE-AC02-06CH11357, and within the frameworks of Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond and the Topical Collaboration in Nuclear Theory Quark Gluon Tomography (QGT) with Award DE-SC0023646.
@@
Our calculations were performed using the Grid Boyle et al. (2016); Yamaguchi et al. (2022) and GPT Lehner et al. software packages. We thank Christoph Lehner for his advice on using GPT. We thank the RBC and UKQCD collaborations for sharing the gauge field configurations used in this work.
@@ Appendix A Bare matrix elements of CG quasi-TMD beam functions in position-space
[…]
@@
Finally, we compare the bare matrix elements extracted from the one-state fit ($\mathrm{Fit}_{\mathrm{st1}}$), two-state fit ($\mathrm{Fit}_{\mathrm{st2}}$), and summation method (SumFit), shown as blue, grey, and purple bands in Fig. 6, respectively. At large time separations, some data points suffer from large statistical uncertainties, limiting their ability to constrain the two-state fit and leading to increased fitting errors. In contrast, results from the summation and one-state fits are more stable. Overall, all three methods yield consistent results, reinforcing the conclusion that the extracted matrix elements are dominated by the ground state, though statistical uncertainties remain notable. As discussed in the main text, we conservatively average the results from the three methods on a sample-by-sample basis for subsequent analysis.
@@ Appendix B CG quasi-TMD beam functions in momentum-space
In the main text, we have presented the renormalized quasi-TMD beam function matrix elements. To compute $\tilde{f}(x,b_{T},P_{z})$, we combine a discrete Fourier transform (DFT) over the region $b_{z}\in[0,b_{z}^{L}]$ with an analytical continuation for the extrapolated region $b_{z}\in[b_{z}^{L},\infty]$. The upper panel of Fig. 10 shows examples of the resulting $x$-dependent quasi-TMD beam functions at various values of $b_{T}$. From left to right, we present the iso-scalar unpolarized, iso-vector unpolarized, and iso-vector helicity cases. As seen in the figure, the signals remain stable across a wide range of $b_{T}$, with no significant decay even at large $b_{T}$, thanks to the CG approach. This is the key that allows us to obtain reliable results up to $b_{T}\approx 1$ fm. We note that these quasi-TMD beam functions are not directly physical, as neither the soft function nor the perturbative matching (as outlined in Eq. (2) of the main text) has been applied at this stage.
In the lower panel of Fig. 10, we compare results obtained with different choices of $b_{z}^{L}$, as well as with a naive DFT performed without extrapolation. The results from different $b_{z}^{L}$ values show excellent agreement, and the differences between the extrapolated and non-extrapolated (DFT-only) cases are minor, although the extrapolated results appear somewhat smoother. This mild difference arises because the matrix elements have already decayed to near-zero beyond $b_{z}\gtrsim 1$ fm. As discussed in the main text, this ensures that the effect of extrapolation—and the associated model dependence—is minimal.
Figure 10: The $x$-dependent quasi-TMD beam functions, obtained from the Fourier transform of renormalized quasi-TMD beam function matrix elements with extrapolation over $b_{z}\in[b_{z}^{L},\infty]$, are shown. The upper panels display results at various values of $b_{T}$ with $b_{z}^{L}=16a$, while the lower panels present results at fixed $b_{T}=8a$ with $b_{z}^{L}$ varying from $10a$ to $16a$. From left to right, examples are shown for the iso-scalar unpolarized, iso-vector unpolarized, and iso-vector helicity cases, respectively.
@@ Appendix C Gauge-invariant quasi-TMDPDFs
The GI quasi-TMD beam functions share the same quark propagators as the CG case but require staple-shaped Wilson lines. Therefore, we also computed the GI quasi-TMD beam functions during the contraction, enabling comparisons between GI and CG approaches. We calculated GI quasi-TMD beam function matrix elements with length of the legs of the stable-shaped Wilson link $\eta=12a$, following the same setup as in Ref. Bollweg et al. (2024). We employed Wilson flow Lüscher (2010), with a flow time $t_{F}=2.0$ to suppress the ultraviolet (UV) fluctuations and improve signal quality.

### v3

@@ Abstract
The bare matrix elements require renormalization. For CG quasi-TMD beam functions, this reduces to a multiplicative constant independent of the quark separation $\{b_{T},b_{z}\}$ and external hadron states [65, 82]. Following Refs. [67, 60], we apply the ratio
$\tilde{h}_{i}^{R}(b_{T},b_{z},P_{z})=\mathcal{N}\frac{\tilde{h}_{i}^{B}(b_{T},b_{z},P_{z},a)}{\tilde{h}^{B,{u_{v}-d_{v}}}_{f_{1}}(b_{T},0,0,a)},$ (7)
where the numerator corresponds to any operator or flavor channel considered in this Letter, and the denominator is fixed to the isovector unpolarized matrix element to cancel the common ultraviolet divergence [65]. $\mathcal{N}=\tilde{h}^{B,u_{v}-d_{v}}_{f_{1}}(0,0,0,a)/\tilde{h}_{i}^{B}(0,0,P_{z},a)$ is a normalization factor. Although this renormalization depends on $b_{T}$, it cancels in the ratios of quasi-TMD beam functions, which we present as the final results. Nevertheless, its inclusion is useful for reducing common systematic errors in intermediate steps such as discretization artifacts and excited-state contamination. This work focuses exclusively on the real part of the matrix element, which provides access to the valence-quark contribution $q_{v}$ in the unpolarized case and the combined quark–antiquark contribution $\Delta q_{+}$ in the helicity case. Previous studies [83, 84, 85] have shown that PDFs (or their even moments) extracted from the real part of the matrix elements agree well with global fits.
Due to confinement, the spatial correlators decay exponentially to zero at large $b_{z}$ [86], making the FT numerically stable. To ensure smooth behavior and suppress the noise at large $b_{z}$, we apply an exponential extrapolation for $b_{z}\gtrsim 1$ fm [87]. The associated systematic uncertainty is estimated by varying the extrapolation range, which has negligible impact on the moderate-$x$ region. After renormalization and FT, we obtain the helicity quasi-TMD beam function $\tilde{g}_{1L}^{\Delta u_{+}-\Delta d_{+}}(x,b_{T},P_{z})$ and the unpolarized quasi-TMD beam functions $\tilde{f}_{1}^{u_{v}-d_{v}}(x,b_{T},P_{z})$ and $\tilde{f}_{1}^{u_{v}+d_{v}}(x,b_{T},P_{z})$, shown in Fig. 2. According to Eq. (2), the quasi-TMD beam function approximates the TMDPDF in $x$-space at $\zeta=(2xP_{z})^{2}$, up to a $b_{T}$-dependent normalization factor—the soft function—as well as matching and power corrections. Since the renormalization factor in Eq. (7) introduces $b_{T}$ dependence, the renormalized quasi-TMD beam function is not monotonic in $b_{T}$. Nevertheless, we observe that it becomes increasingly suppressed as $b_{T}$ grows when $x\to 1$, indicating that the $O(1/(xP_{z}b_{T}))$ power correction is suppressed, which is consistent with previous lattice findings [58, 59, 59, 67, 69]. The statistical errors of the quasi-TMD beam function decrease with increasing $x$ and become small in the valence quark region. Although these errors grow with increasing $b_{T}$ due to the worsening signal-to-noise ratio, they remain sufficiently controlled to allow for meaningful predictions. Next, we construct the RGI ratios of the quasi-TMD beam functions. As a cross-check, we also compute the gauge-invariant (GI) quasi-TMD beam functions using the same lattice setup, and confirm that their RGI ratios are consistent with those from the CG method, as expected from Eq. (2[43, 47, 66]; see Appendix C.
@@
Figure 3: Lattice QCD results (filled symbols with error bars) for the ratio $R^{u-d}_{g_{1L}/f_{1}}(x,b_{T})$ of isovector helicity to isovector unpolarized TMDPDFs in Eq. (3), shown as a function of $b_{T}$ at several values of $x$. The hatched bands indicate the corresponding global fits from MAP-Heli [33] and TNTC [32].
@@
Figure 4: Lattice QCD results (filled symbols with error bars) for the ratio $R_{f_{1}}^{u/d}(x,b_{T})$ of up- to down-quark unpolarized TMDPDFs in Eq. (4), shown as a function of $b_{T}$ at several values of $x$. The hatched bands denote the corresponding global fits from MAP24 [31] and ART25 [29].
Figure 3 shows the RGI ratio $R^{u-d}_{g_{1L}/f_{1}}(x,b_{T})$, defined in Eq. (3), for several values of $x$. Across all $x$ in the range $[0.2,0.6]$, the ratio is consistent with being constant in $b_{T}$ within errors. This indicates that within the Collins-Soper-Sterman (CSS) [3] formalism of parametrizing the TMDPDFs, the intrinsic nonperturbative TMD part is not sensitive to spin. At smaller $x$, particularly around $x=0.2$, our results are consistent with the MAP-Heli fit [33], while some deviations emerge at larger $x$. However, the TNTC fit [32] yields substantially lower values across the entire $x$ range, although its $b_{T}$-dependence is also mild within uncertainties. The observed discrepancies between MAP-Heli and TNTC could be due to the experimental precision, analysis method and model uncertainties. While our lattice calculation is performed at a single lattice spacing and momentum, and thus subject to discretization effects and power corrections, these effects have been shown to be under control in similar LaMET studies, such as the pion valence quark PDF [88], the CS kernel [67], and the proton unpolarized PDF [84]. Therefore, our lattice results in the moderate-$x$ region can meaningfully differentiate and constrain phenomenological parametrizations of the TMDPDFs.
Figure 4 shows the ratio $R_{f_{1}}^{u/d}(x,b_{T})$ of the valence $u$ to $d$ unpolarized TMDPDFs, as defined in Eq. (4), for several values of $x$. For comparison, results from the flavor-dependent global fits MAP24 [31] and ART25 [29] are shown as bands. Better agreement in the trends in $b_{T}$ is observed between lattice and ART25, which also show a flat $b_{T}$-dependence across all $x$. Nevertheless, the magnitudes of the two results are noticeably different, which is subject to their own systematic uncertainties. At small $b_{T}$, some similarity between the lattice results and MAP24 emerges, likely due to the dominance of input collinear PDFs in their analysis. The mild $b_{T}$ variation in our lattice results indicates that the intrinsic nonperturbative TMD part of CSS parameterization has a weak dependence on the quark flavor. Notably, while the $b_{T}$-dependence appears flat, the ratio still shows a noticeable dependence on $x$, indicating nontrivial longitudinal structure. Our finding is particularly valuable in light of the tensions among existing global fits, where incorporating flavor dependence introduces additional parameters that may increase statistical instability and contribute to systematic uncertainty. Therefore, our lattice QCD results offer a first-principles benchmark to guide future global analyses and help reduce the associated uncertainties.
Conclusion. We presented the first lattice QCD calculation of the isovector helicity TMDPDF and the flavor-dependent unpolarized TMDPDFs for up and down quarks in the proton. This calculation used domain wall fermion configurations with physical quark masses, and the CG method was employed to reliably access the deep nonperturbative region up to $b_{T}\sim 1$ fm.
We obtain ratios of the TMDPDFs, which can be directly compared to those from global analyses. Our results show a close similarity between the helicity and unpolarized isovector TMDPDFs at moderate $x$, as well as a mild flavor dependence in the unpolarized case. With the current precision, our results can differentiate phenomenological parametrizations of the TMDPDFs and be used as valuable inputs to constrain future global analyses.
The present study provides first benchmarks at moderate $x$ on a single lattice spacing and volume, and neglects disconnected diagrams in the unpolarized channel. Future work will extend these results by increasing the proton momentum, employing finer lattices, and addressing additional systematics—including excited-state contamination, disconnected diagrams, and finite-volume effects—so as to reach precision-level control. In parallel, together with a lattice determination of the soft function [41, 69], these improvements will enable precise predictions of the proton’s three-dimensional structure with full spin dependence.
@@
Acknowledgements. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704, Contract No. DE-AC02-06CH11357, and within the frameworks of Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond and the Topical Collaboration in Nuclear Theory Quark Gluon Tomography (QGT) with Award DE-SC0023646.
@@
Our calculations were performed using the Grid [89, 90] and GPT [91] software packages. We thank Christoph Lehner for his advice on using GPT. We thank the RBC and UKQCD collaborations for sharing the gauge field configurations used in our work.
@@ Appendix A Bare matrix elements of CG quasi-TMD beam functions in position-space
$\begin{aligned} C^{\rm 3pt}_{\mathcal{P},\Gamma}(t_{s},\tau;b_{T},b_{z},P_{z})=\sum_{m,n}A_{m}^{*}A_{n}\langle m|O|n\rangle e^{-\tau E_{n}}e^{-(t_{s}-\tau)E_{m}},\end{aligned}$ (11)
@@
[…]
@@ Appendix B CG quasi-TMD beam functions in momentum-space
At large distances, spatial correlators are known to decay exponentially [86]. As shown in the upper panel of Fig. 11, the renormalized matrix elements approach zero for $b_{z}\gtrsim 1$ fm, within the uncertainties. To perform the FT over $b_{z}$ and extract the quasi-TMDs in $x$-space, we model the long-distance behavior by fitting the data in the range $b_{z}\in[b_{z}^{L},b_{z}^{L}+4a]$ using an exponential function, $Ae^{-mb_{z}}$, with a prior constraint of $m>0.1$ GeV. In the upper panel of Fig. 11, we illustrate the fit results for $b_{z}^{L}=16a$ as the bands, which effectively capture the data behavior, although the uncertainties decrease more rapidly than the data itself. To account for potential systematic uncertainties arising from the extrapolation, we vary the fitting range by adjusting $b_{z}^{L}$ between $10a$ and $16a$. Finally, to compute $\tilde{f}(x,b_{T},P_{z})$, we combine a discretized FT for the range $b_{z}\in[0,b_{z}^{L}]$ with an analytical FT for the extrapolated region $b_{z}\in[b_{z}^{L},\infty]$. This hybrid approach ensures a smooth and reliable reconstruction of the quasi-TMDs in momentum space.
In the lower panel of Fig. 11, we compare results obtained with different choices of $b_{z}^{L}$, as well as with a naive DFT performed without extrapolation. The results from different $b_{z}^{L}$ values show excellent agreement, and the differences between the extrapolated and non-extrapolated (DFT-only) cases are minor, although the extrapolated results appear somewhat smoother. This mild difference arises because the matrix elements have already decayed to near-zero beyond $b_{z}\gtrsim 1$ fm. As discussed in the main text, this ensures that the effect of extrapolation and the associated model dependence is minimal.
Figure 11: Upper panel: Real parts of the renormalized matrix elements for Coulomb-gauge (CG) quasi-TMD beam functions at nucleon momentum $P_{z}=1.62$ GeV, plotted as functions of $b_{z}$ for several values of $b_{T}$. The bands indicate exponential fits used for extrapolation beyond $b_{z}^{L}=16a$. Lower panel: $x$-dependent quasi-TMD beam functions at fixed $b_{T}=8a$, obtained via discrete FT (DFT) with and without exponential extrapolation. Results are shown for $b_{z}^{L}$ ranging from $10a$ to $16a$. From left to right, examples are shown for the iso-scalar unpolarized, iso-vector unpolarized, and iso-vector helicity cases, respectively.
@@ Appendix C Gauge-invariant quasi-TMDPDFs
The GI quasi-TMD beam functions share the same quark propagators as the CG case but require staple-shaped Wilson lines. Therefore, we also computed the GI quasi-TMD beam functions during the contraction, enabling comparisons between GI and CG approaches. We calculated GI quasi-TMD beam function matrix elements with length of the legs of the stable-shaped Wilson link $\eta=12a$, following the same setup as in Ref. [67]. We employed Wilson flow [92], with a flow time $t_{F}=2.0$ to suppress the ultraviolet (UV) fluctuations and improve signal quality.